OpenStudy (domebotnos):

MEDAL AND FAN Solve the system of equations. x + 3y = -1 2x + 2y = 6 (-4, 1) (2, -1) (4, -1) (5, -2)

OpenStudy (samanthagreer):

I'll use elimination method. 3x + 8y = 12 2x + 2y = 3 -6x -16y = -24 6x + 6y = 9 -10y = -15 /-10 /-10 y = 1.5 Sub in y into the equation and solve for x http://www.google.com/url?sa=t&rct=j&q=&esrc=s&source=web&cd=1&ved=0CCMQFjAA&url=http%3A%2F%2Fopenstudy.com%2Fupdates%2F52b49b19e4b03e81ff961efc&ei=-s_aVMbiB_jasASzhoG4Aw&usg=AFQjCNFRL5A88vVRFBO5-u3dAI6lCZnNLw&bvm=bv.85761416,d.cWc

OpenStudy (sleepyjess):

You can also just substitute in the points...

OpenStudy (anonymous):

DOMEBOTNOS USE THE SUBSTITUTION METHOD! HERE. MAKE X+3y=-1 and solve for X so you can SUBSTITUTE THAT INTO THE X OF THE SECOND EQUATION So WE GET X=-3y-1, PLUG THAT INTO THE OTHER EQUATION 2(X)+2Y=6, PLUG IN -3y-1 here 2(-3y-1)+2y=6 -6y-2+2y=6... Solve for Y and then plug the value of Y into any of the two original equations and you will get an X value, Those are your two answers which can be written in ordered pair (x,y)

OpenStudy (domebotnos):

THANK YOU GUYS SO MUCH :_))))